For the formation of 3.65 gram of HCL gas, what volume of hydrogen gas and chlorine gas are required at NTP condition ? (1) 1L,1L (2) 1.12L , 2.24L (3) 3.65 L, 1.83L (4) 1.12L, 1.12L

Grade:12th pass

5 Answers

Saurabh Koranglekar
4 years ago
Dear student

At mass of HCL is 36.5

And reaction is H2 + Cl2 = 2 HCL

0.1 mole of HCl req 0.05 moles of reactants

0.05 *22.4 = 1.12 lit

Regards
Vikas TU
4 years ago
Moles of HCl = given mass / molar mass  = 3.65 /36.5 = 0.1 mole 
0.5H2 + 0.5 Cl2 = HCl 
From the given balanced equation:
1 mole of HCl require = 0.5 mole of Hydrogen and 0.5  mole of chlorine
0.01 mole of HCl require=0.05 mole of Hydrogen and 0.05 mole of chlorine.
According to Avogadro's law, 1 mole of gas occupies=22.4 L of gas at STP.
Thus 0.05 mole of every gas occupies= Litres at STP.
 
Vikas TU
4 years ago
Dear student 
the ans would be 1.12 Litre at STP , 
Hope you can solve and understand the method ,
Good luck cheers
Om Biradar
4 years ago
As the Ratio of H to Cl in HCl is 1:1
we know that at a constant temperature the volume of equal moles of two gases is equal
Volume of H2 gas = volume of Cl2 gas
This is true only for option A and D 
Therefore options C and D can be neglected 
 
By avagadro ,
molar mass of 1 mole of HCl = molar mass of H + molar mass = 1+35=36g
We know that 36g or 1mole of HCl occupies 1 liter volume at STP 
=> 3.68 grams would hold 3.68/36 = 1.24l
Therefore we would require equal amounts of H and Cl to make HCl 
Volume of H = Volume of Cl = Volume of HCl /2 = 0.1 /2 = 0.05l 
Om Biradar
4 years ago
This is the continuation of the before and 
Volume of H and Cl = 1.24/2 l = 0.62 l 
We want the volume of H2 gas 
We know that the volume of a diatomic gas is half of that of its monoatomic gas 
=> volume H2 gas and Cl2 gas which is required to make 3.68g of HCl 
= 0.62 /2 = 0.31L 
 Thus the correct amns is none of the above (which is not an option) 
Therefore you van conclude that the question is incorrect